OpenStudy (anonymous):

Use substitution 1. y=x+6 y=2x a.(6,12) b.(-12,-6) c.(-6,-12) d.(2,4)

terenzreignz (terenzreignz):

y = x + 6 y = 2x since both of them are y, then they must be equal to each other, right?

OpenStudy (anonymous):

im not really sure

terenzreignz (terenzreignz):

x + 6 is equal to y 2x is equal to y so they must be equal to each other, since they both equal y. x + 6 = 2x So just solve for x ^_^

OpenStudy (anonymous):

A?

terenzreignz (terenzreignz):

That is correct :)

OpenStudy (anonymous):

what about this one 2. 3x+2y=7 -3+11=y a.(6,-3 b.(6,-7) c(-4,9/12) d.(5,-4)

terenzreignz (terenzreignz):

For one thing, I'm sure you missed something in that second equation. And... is this a quiz?

OpenStudy (anonymous):

no thats the equation and no its not

OpenStudy (anonymous):

-3x

terenzreignz (terenzreignz):

I see. Well what you do here is replace the y in the first equation with -3x + 11 and then solve for x.

OpenStudy (anonymous):

how do i solve for x?

terenzreignz (terenzreignz):

First, replace the y in the first equation with -3x + 11. Show me what you get.

OpenStudy (anonymous):

im confused

terenzreignz (terenzreignz):

Okay, how about I present an example?

OpenStudy (anonymous):

it might help

terenzreignz (terenzreignz):

Will it though? ^_^ Suppose we have 2x + 8y = 22 y = x - 1 so, you see that y in the first equation? We replace it with x - 1, which comes from the second equation. Like so: 2x + 8(x - 1) = 22 Now do likewise

OpenStudy (anonymous):

so like -3x+11(x-1)=7?